help p[lzzzExercise 5A-3 Cost Behavior; High-Low Method [LO5-10] Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singa
Answer #1

a.) As per high low method

Variable cost = difference in cost/ difference in untis(kilometer)

= 2,501,200-2,053,200/174,000-116,000

= 448,000/58,000

= 7.724 variable cost per kilometer

Calculation of fixed cost

Consider units of 174,000 kilometers

Variable cost = 174,000*7.724 = 1,344,000

Therefore fixed cost = 2,501,200-1,344,000 = 1,157,200

b.) Given total cost for driving a truck for 174,000 kilometers is

2,501,200

Therefore fixed cost + 174,000*variable cost per unit = total cost

I.e, 1,157,200+(174,000*7.724)

= 1,157,200+1,344,00

= 2,501,200

c.) Calculation of operating cost for 145,000 kilometers

= Fixed cost + variable cost

= 1,157,200+ (145,000*7.724)

= 1,157,200+1,120,000

= 2,277,200